OpenStudy (anonymous):

*C=1.25N+250 C=cost =# of tickets sold *T=4N T=total amount collected N=#of tickets a. what is the cost if 213 tickets are sold? b. how much are the total ticket sales if 213 tickets are sold? c. what is the drama clubs profit or loss if 213 tickets are sold? d. if the total ticket sales are $780, how many people attended? e. what is the cost of putting on the play for the number of people you found in part (d)?

OpenStudy (anonymous):

what grade of math are u in?

OpenStudy (anonymous):

7th

OpenStudy (anonymous):

y

OpenStudy (anonymous):

A. you would put in 213 for the cost and subtract 2.50 from each side then divide by 1.25

OpenStudy (anonymous):

so 168.4?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

ok wat about the others?

OpenStudy (anonymous):

okay so B. you would multiply 1.25 and 213 and hen add 2.50

OpenStudy (anonymous):

so,268.25?

OpenStudy (anonymous):

yes.. thats what i got.

OpenStudy (anonymous):

ok and c

OpenStudy (anonymous):

okay C. i don't get. are u sure u spelled everything right?

OpenStudy (anonymous):

yes i spelled everything right

OpenStudy (anonymous):

Is this the equation ? C = 1.25N + 250 or is it 2.50 ? C = cost N = number of tickets Wouldn't the total amount collected be the same as the cost ? The T = 4N is confusing me. Are the tickets 1.25 each or 4.00 each ?

OpenStudy (anonymous):

the tickets r worth 4.00

OpenStudy (anonymous):

250 and the tickets are 4.00 but the theater is charging the drama club 1.25 for each ticket sold

OpenStudy (anonymous):

I am having trouble understanding this question

OpenStudy (anonymous):

yea and y did u delete it?

OpenStudy (anonymous):

Because I am confused......with the T and the C ?

OpenStudy (anonymous):

C is the total cost for the drama club to put on the play and T is the total amount collected from ticket sales

OpenStudy (anonymous):

Let me try again..... (a) C = 1.25N + 250 C = 1.25(213) + 250 C = 266.25 + 250 C = 516.25 (b) T = 4N T = 4(213) T = 852 (c) Drama's club profit is 852 - 516.25 = 335.75 (d) T = 4N 780 = 4N 780/4 = N 195 = N (e) C = 1.25N + 250 C = 1.25(195) + 250 C = 243.75 + 250 C = 493.75 I am still a little confused, so you might want to get a second opinion on this

OpenStudy (anonymous):

@Opcode ....this is really messing with my brain...help

OpenStudy (anonymous):

@Mertsj ....can you take a peek at this....I am confused

OpenStudy (anonymous):

This should be easy and I don't know why it is confusing me :(

OpenStudy (opcode):

A is correct that I'm sure of.

OpenStudy (anonymous):

On B are the tickets $4 or $1.25 ?

OpenStudy (opcode):

I agree with you on B. Because it says: *T=4N T=total amount collected N=#of tickets

OpenStudy (anonymous):

1.25

OpenStudy (anonymous):

i mean 4

OpenStudy (mertsj):

B. Total=4n If 213 tickets are sold, total collected is 4 x 213

OpenStudy (mertsj):

C. to get profit or loss subtract cost from sales. In other words, subtract the answer to a from the answer to b

OpenStudy (mertsj):

D. T = 4n If T = 780 replace T with 780 and solve that equation

OpenStudy (mertsj):

E. Take the answer to d and put it into the cost equation: C=1.25n+250

OpenStudy (anonymous):

thank you so much @Mertsj ...your a lifesaver

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

We got it @Giant_Cupcake ........I don't know why I was so confused on that

OpenStudy (anonymous):

yes but on e its 493.75?

OpenStudy (mertsj):

yes

OpenStudy (anonymous):

ok thank you so much

OpenStudy (mertsj):

yw

OpenStudy (anonymous):

Let me check again.... C = 1.25N + 250 C = 1.25(195) + 250 C = 243.75 + 250 = 493.75 thats what I got :)

OpenStudy (anonymous):

@Mertsj is the best....so is @Opcode

OpenStudy (anonymous):

theres a f and a g will u guys help

OpenStudy (anonymous):

I will give it a try

OpenStudy (anonymous):

ok. f. how many tickets does the drama club need to sell to break even? g. the drama club would like to earn a profit of $500 from the play. how many tickets need to be sold for the club to meet this goal?

OpenStudy (anonymous):

Lets figure this out......Is it only costing the drama club $ 250 ?

OpenStudy (anonymous):

no props cost $250 and the theater is charging $1.25 for each ticket sold and each ticket costs $4.00

OpenStudy (anonymous):

1.25N + 250 = 4N 250 = 4N - 1.25N 250 = 2.75N 250/2.75 = 2.75N/2.75 90.91 = N (round to 91) check... 1.25(91) + 250 = 4(91) 113.75 + 250 = 364 363.75 = 364 I am not sure...I must be doing something wrong because I am a quarter off. You are probably going to have to ask someone else....I am sorry

OpenStudy (anonymous):

so ur answer for f is ~91 and for g?

OpenStudy (anonymous):

I am not sure....I don't seem to be getting the answer....you are probably going to have to ask someone else....again I am sorry

OpenStudy (anonymous):

its ok thank you for the help and by the way while u were trying to figure out all of my problems i was doing other ones lol

OpenStudy (anonymous):

nothing wrong with that.....if they got to be done, you must do them when you can :)

OpenStudy (anonymous):

lol ikr and im finally done!!!!!!(:

OpenStudy (anonymous):

and im going to get my brother to look at them

OpenStudy (anonymous):

Got to go now.....see ya later @Giant_Cupcake ......get your brother to check them.....good luck

OpenStudy (anonymous):

i will ok goodnight smart person

OpenStudy (anonymous):

you are smart as well....don't ever think that you are not :)

OpenStudy (anonymous):

ik im smart im just slow

OpenStudy (anonymous):

Good night now....I am gonna go dream about some sexy guys...lol

OpenStudy (anonymous):

lol wow im gonna dream about my sexy bf lol